The energy transfer to a device is 112 000 J when there is a potential difference of 185 volts across it. If the current through the device is 0.0644 A, how long in hours is the current flowing? 


1
Expert's answer
2021-12-20T10:27:19-0500

The power supplied by the device is given by

"P=V I"

Where.

"V=185\;V" Is the voltage.

"I=0.0644 A" Is the current.


Evaluating numerically.

"P=V I\\
P=185\;V\times 0.0644\;A\\
P=11.9\;W"

But remember that the power is given by

"P=\dfrac{E}{t}"

Where.

"E=112 000\;J" Is the energy.

"t" Is the time.


Obtaining the expression for time

"P=\dfrac{E}{t}\\
P\;t=E\\
t=\dfrac{E}{P}\\"


Evaluating numerically in seconds.

"t=\dfrac{E}{P}\\
t=\dfrac{112 000 J }{11.9\;W}\\
t=9411.8\;s"

The time in which the current flows is.

"t=9411.8\;s"